OpenStudy (anonymous):

In 1995 the life expectancy of a male in a certain country was 63.6 years In 1999 it was 67.4 years Let E represent the life expectancy in year t and let t represent the number of years since 1995 The linear function E(t) that fits the data is E(t)= t + (round to the nearest tenth) Use the function to predict the life expectancy of males in 2006 E(11)= (round to the nearest tenth)

OpenStudy (dumbcow):

E(0) = 63.6 E(4) = 67.4 --> (67.4 - 63.6)/4 = 3.8/4 = 0.95 E(t) = 0.95t + 63.6 --> E(11) = 0.95(11) + 63.6 = 74.05